OpenStudy (sonyalee77):

PLEZZZZZZ HELP!!!!!!!!!!! Solve and tell steps please : 3y-2x=7

OpenStudy (sonyalee77):

Any one plez I really need this

OpenStudy (anonymous):

there should be another equation

OpenStudy (sonyalee77):

the real qestion was this but i really wanted to do it step by step: The solution of 3y - 2x = 7, with x = 4, is y = 4. true or false

OpenStudy (anonymous):

well 3(4) - 2(4) = 12 - 8 4 4 isn't 7 so that would have to be false

OpenStudy (sonyalee77):

thank you